-8z=4x
3x+z=2y
y-z=2x-1 solve the matrices

Answer by Alan3354(69443)

-8z=4x --> x = -2z
3x+z=2y
y-z=2x-1
-------------
3x+z=2y
Sub for x
--
3(-2z) + z = 2y
y = -5z/2
----------
y-z=2x-1
Sub for x and y
-5z/2 - z = 2(2z) - 1
-7z/2 = 4z - 1
-15z/2 = -1
z = 2/15 *************
y = -1/3 *************
x = -4/15
====================